### v1

# Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registrations

Abstract
In this paper, we discuss Mahalanobis k-NN: a statistical lens designed to address the challenges of feature matching in learning-based point cloud registration when confronted with an arbitrary density of point clouds, either in the source or target point cloud. We tackle this by adopting Mahalanobis k-NN’s inherent property to capture the distribution of the local neighborhood and surficial geometry. Our method can be seamlessly integrated into any local-graph-based point cloud analysis method. In this paper, we focus on two distinct methodologies: Deep Closest Point (DCP) and Deep Universal Manifold Embedding (DeepUME). Our extensive benchmarking on the ModelNet40 and Faust datasets highlights the efficacy of the proposed method in point cloud registration tasks. Moreover, we establish for the first time that the features acquired through point cloud registration inherently can possess discriminative capabilities. This is evident by a substantial improvement of about 20% in the average accuracy observed in the point cloud few-shot classification task benchmarked on ModelNet40 and ScanObjectNN. The code is publicly available at https://github.com/TejasAnvekar/Mahalanobis-k-NN
Keywords:
Point Cloud Registration Metric Distance 3D Vision.
Figure 1: The visual supremacy of our proposed methodologies, MDCP-v1 and MDeepUME, becomes apparent in a point cloud registration task involving a target point cloud with only half the points compared to the source point cloud. In the visualization, highlighted boxes in red illustrate the limited performance of DCP-v1 [DCP] and DeepUME [DeepUME], while the highlighted boxes in green demonstrate the resilience of the proposed Mahalanobis versions of DCP-v1 [DCP] and DeepUME [DeepUME].
1 Introduction
With the rapid progress of high-precision sensors like LiDAR [lidar] and Kinect [kinect], point clouds have become the prevalent data format for 3D representation. Due to sensor limitations in capturing only partial views, registration algorithms are crucial for amalgamating scans into comprehensive 3D scenes. Point cloud registration involves determining transformation matrices between point cloud pairs, facilitating the fusion of partial scans to produce coherent 3D representations.
Registration is a fundamental task spanning various computational fields such as medical imaging [medicalImaging], robotics [robotics], autonomous driving [autoDriving], and computational chemistry [ComChe]. Estimating transformation matrices serves applications like 3D reconstruction [reconstruction] and precise 3D localization [Localization]. It underpins the development of intricate 3D maps for autonomous driving [autoDriving], environment reconstruction in robotics [robotics], and improved safety in mining. Additionally, registration’s ability to facilitate high-precision localization proves invaluable for entities like driverless cars [autoDriving], ensuring precise positioning and interaction within the 3D environment.
The registration of point clouds has garnered significant attention through closed-form [ICP, goICP, FGR, aligningPointCloudView, ume] and learning-based [pointnetLK, DCP, DeepGMR, DeepUME, rgm] methodologies. DCP [DCP] tackles the challenge of feature matching (correspondences) by initially estimating local features via weight-shared DGCNN [dgcnn] (Dynamic Graph Convolutional Neural Network). Subsequent steps involve an attention-based module and differentiable SVD modules for point-to-point registration. Conversely, DeepUME [DeepUME] adopts a distinct approach by projecting points into a space invariant under $\mathcal{SO}(3)$ transformations. This space is leveraged to compute invariant per-point features that address real-world data sampling issues. Employing two weight-shared transformers and DGCNN [dgcnn], correspondences are established and fed into UME [ume] for rigid transformation parameter estimation. A limitation of both DCP [DCP] and DeepUME [DeepUME] is susceptibility to failure in scenarios where either the source or target point cloud exhibits arbitrary (lower) density than the others—an occurrence commonly encountered in 3D scan registration from diverse sensors. This limitation is attributed to the methods’ reliance on estimating per-point correspondences using local features extracted from edge-conv [dgcnn], which operates on graphs constructed using k-nearest neighbors (k-NN).
We introduce Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration to mitigate this issue. The efficacy of Mahalanobis distance as an evaluation metric has already been demonstrated in reference-based point cloud quality assessment [MD1]. Our proposed Mahalanobis k-NN can be used as a plugin for any point-cloud registration method. In this paper, we propose mahalnobis versions of DCP-(v1,v2) [DCP] and DeepUME [DeepUME]. Both methods operate on diverse point-cloud representation spaces. This augmentation facilitates pointcloud registration as shown in Figure 1; across various publicly available 3D datasets, encompassing diverse benchmarking scenarios, including: 1) unseen-category evaluation, 2) robustness to various noise types, and 3) efficiency towards varying point densities. Additionally, we demonstrate the discriminative capabilities of the proposed Mahalanobis lens through point cloud few-shot classification tasks. In this context, models pre-trained for registration tasks leverage DGCNN [dgcnn] features for few-shot evaluation on both ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets. Finally, we summarize our contributions as follows:
  • We propose Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registration by incorporating Mahalanobis distance in:
    • Deep Closest Point (DCP) [DCP] that operates on Euclidean coordinate space.
    • Deep Universal Manifold Embedding (DeepUME) [DeepUME] that operates on $\mathcal{SO}(3)$ invarient space.
  • We demonstrate the efficacy of the proposed method on various point cloud registration tasks on a variety of publicly available data sets and compare the results with state-of-the-art techniques.
  • We perform an endurance test to evaluate the robustness and generalization of the proposed method. We achieve state-of-the-art results compared to other point cloud registration methods on all benchmarking strategies and on almost all evaluation metrics.
  • To the best of our knowledge, we are the first to demonstrate discriminative prowess inherited in point cloud registration task by benchmarking point cloud few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] datasets through features extracted by DCP [DCP], we report up to an average of 20% increment when Mahalanobis is incorporated.
  • We will release our code to facilitate reproducibility and future research upon acceptance.
2 Related Works
Optimization Based. Preceding the advent of the deep learning era in 3D point cloud registration, one common strategy involved extracting and matching spatially local features, as demonstrated in studies like [comprehensiveEvalFeatureDescriptor, spinImages, aligningPointCloudView, RepeatabilityRobust3Dmatching, fastRobustDescriptor, rotationalContour]. Many existing methods in this category are adaptations of 2D image processing solutions, such as variants of 3D-SIFT [3Dsift] and the 3D Harris key-point detector [harris3D]. However, key-point matching in 3D presents challenges due to the absence of a regular sampling grid, artifacts, and sampling noise, leading to high outlier rates and localization errors. To achieve global alignment, key-point matching typically utilizes outlier rejection methods like RANSAC [ransac] and is refined using local optimization algorithms [ICP, scanRegistration3Dndt, registrationWithoutICP, iterativePointMatching]. Notably, DGR [DGR] follows a similar paradigm, but it incorporates learnable inlier detection. Researchers have proposed numerous works for handling outliers and noise [trimmedICP], formulating robust minimizers [robustRegistration2D3D], and devising more suitable distance metrics. The widely used Iterative Closest Point algorithm (ICP) [ICP, iterativePointMatching], a popular refinement algorithm, constructs point correspondences based on spatial proximity and employs a transformation estimation step. Over time, various ICP variants [efficientICP, goICP, sparseICP] have been proposed to enhance convergence rate, robustness, and accuracy in the 3D point cloud registration.
Learning Based. Registration approaches leverage learning-based techniques alongside closed-form methods. Pioneered by PointNet [pointnet] and advanced by DGCNN [dgcnn], learning from data for point cloud registration has become powerful (e.g., PointNet++, DCP [DCP], PRNet [prnet]). PointNetLK [pointnetLK] minimizes learned feature distance using a differentiable Lucas-Kanade algorithm, while DCP uses attention and differentiable SVD for point-to-point registration. Recent work [rgm] introduces graph transformation and deep graph matching. Moreover, DeepUME [DeepUME] employ $\mathcal{SO}(3)$ invariant coordinate systems for global registration. DeepGMR [DeepGMR] tackles pose-invariant correspondences between raw point clouds and Gaussian-mixture models but suffers from noise sensitivity. DeepUME [DeepUME] fuses closed-form Universal Manifold Embedding (UME) and deep learning to handle sparse and unevenly sampled point clouds with large transformations. This framework is trained end-to-end and learns joint-resampling strategies and invariant features.
DCP [DCP] and DeepUME [DeepUME] achieve impressive performance but struggle when the source or target point cloud densities are unequal. Both rely on DGCNN [dgcnn] for local feature-based correspondence estimation. However, DGCNN’s Euclidean metric for k-nearest neighbors lacks surface awareness. We propose Mahalanobis versions of DCP and DeepUME, introducing the Mahalanobis distance for surface-aware features and enhanced registration in real-world scenarios. Furthermore, our proposed approach demonstrates remarkable discriminative capabilities, as evident from its performance in point cloud few-shot classification tasks.
3 Problem Setting
We denote $x_{i}^{T}(i\in[1,M])$ and $y_{i}^{T}(i\in[1,N])$ as row vectors from matrices $X\in\mathbb{R}^{M\times 3}$ and $Y\in\mathbb{R}^{N\times 3}$ respectively. The matrices $X$ and $Y$ represent two distinct point clouds, often originating from two scans of the same object (we assume $Y$ is transformed from $X$ by an unknown rigid motion). Here, $x_{i}$ and $y_{j}$ represent the coordinates of the $i$th point in $X$ and the $j$th point in $Y$, for $i\in[1,M]$ and $j\in[1,N]$ respectively. Assuming that there are $S$ pairs of corresponding points between the two point clouds, the goal of registration is to determine the optimal rigid transformation parameters $g$ (comprising a rotation matrix $R\in\mathcal{SO}(3)$ and a translation vector $t\in\mathbb{R}^{3}$) that aligns point cloud $X$ to point cloud $Y$, as illustrated below:
$\operatorname*{arg\,min}_{R\in\mathcal{SO}(3),\,t\in\mathbb{R}^{3}}\left\|d\left(X,g(Y)\right)\right\|_{2}^{2}$ (1)
The term $d\big{(}X,g(Y)\big{)}$ is a measure of the projection error between $X$ and the transformed $Y$, expressed as $d\big{(}X,RY+t\big{)}=\sum_{s=1}^{S}\left\|x_{s}-\big{(}Ry_{s}+t\big{)}\right\|_{2}$, where $s$ iterates over the $S$ pairs of corresponding points. The optimization problem in Equation 1 embodies a well-known “chicken-and-egg” scenario: determining the optimal transformation matrix requires knowledge of true correspondences  [correspondance1, correspondance2]; conversely, accurate correspondences can be established given the optimal transformation matrix. However, the simultaneous resolution of both aspects poses a non-trivial challenge.
4 Method
In this paper, we focus on point cloud registration using DCP [DCP] and DeepUME [DeepUME] for global one-shot rigid transformations (contrasting iterative methods like ICP [ICP]). Both rely on DGCNN [dgcnn] for local features but use static k-nearest neighbor (k-NN) graphs, limiting semantic space creation. Unlike, DGCNN’s dynamic k-NN graphs, these methods restrict feature projection depth. We propose using the Mahalanobis distance [MD] to enhance k-NN, enabling surface-aware neighbor selection based on principal components. This extracts surface and corner features crucial for accurate matching and robust transformation estimation.
4.1 Mahalanobis k-NN
Figure 2: The illustration contrasts Euclidean distance fields (red) capturing spatial neighbors with Mahalanobis distance fields (blue) considering the underlying data distribution. A chair point cloud exemplifies this. The black query point is surrounded by Euclidean neighbors (red points) and Mahalanobis neighbors (blue points). The depiction clearly illustrates the impact of Mahalanobis distance, effectively capturing surface points per the data distribution vital for precise feature matching.
In this section, we introduce Mahalanobis distance as a statistical lens for robust feature matching in DCP [DCP] and DeepUME [DeepUME]. We modify DGCNN to construct a graph $\mathcal{G}^{\dagger}$ using Mahalanobis k-nearest neighbors (k-NN) for enhanced feature selection. DGCNN constructs a graph $\mathcal{G}$, applies nonlinearity for edge values, and performs vertex-wise aggregation ($\max$ or $\sum$) in each layer. Let $x_{i}^{l}$ be the embedding of point $i$ in the $l$-th layer, and $h_{\theta}^{l}$ be a nonlinear function parameterized by a shared multi-layer perceptron (MLP). DGCNN can be formulated as:
$x^{l}_{i}=f\Big{(}\big{\{}h_{\theta}^{l}(x_{i}^{l-1},x_{j}^{l-1})\forall j\in\mathcal{N}_{i}\big{\}}\Big{)}$ (2)
where $\mathcal{N}_{i}$ denotes neighbors of vertex $i$ in $\mathcal{G}$.
We propose a novel augmentation by integrating Mahalanobis k-nearest neighbors (k-NN) in Equation.2 to construct graph $\mathcal{G}^{\dagger}$. This graph prioritizes surface awareness through selection of surficial neighbors $\mathcal{N}_{i}^{\dagger}$ based on Mahalanobis distance as shown in Figure 2.
To compute Mahalanobis distance between points $x_{i}$ and $x_{j}$, we use:
$D_{M}(x_{i},x_{j})=\sqrt{(x_{i}-x_{j})^{T}\cdot C^{-1}\cdot(x_{i}-x_{j})}$ (3)
where: $D_{M}(x_{i},x_{j})$ is Mahalanobis distance between $x_{i}$ and $x_{j}$, and $C$ is covariance matrix of data points. One can argue, geodesic graph-based methods capture similar or sometimes better surficial information. To counter this, we implement vectorized Floyd-Warshall (refer supplementary for implementation detail on vectorized Floyd-Warshall) on the Euclidean-knn graph and observe that the performance is slower and also not as effective as the proposed Mahalanobis variants of DCP and DeepUME.
The impact of Mahalanobis distance is visually depicted in Figure 2. This surficial awareness proves crucial in feature matching. Incorporating Mahalanobis distance into DCP  [DCP] and DeepUME  [DeepUME] yields superior performance in point cloud registration. Discriminative power is affirmed in point-cloud few-shot classification, with pre-trained DGCNN models from vanilla DCP and Mahalanobis DCP showcasing remarkable classification capabilities.
5 Experiments
In this section, we showcase the superior performance of our proposed method in the domain of point cloud registration when compared against existing state-of-the-art techniques such as ICP [ICP], GO-ICP [goICP], FGR [FGR], PointNetLK [pointnetLK], DCP [DCP], and DeepUME [DeepUME] on publicly available datasets. Our evaluation encompasses a spectrum of benchmarking strategies, encompassing scenarios where one point cloud is notably sparser than the other. For a fair comparison, we implemented a vectorized Floyd-Warshall algorithm in PyTorch to compute geodesic distances and compare it to our proposed method for surface feature extraction. We further validate the Mahalanobis metric’s discriminative power in few-shot learning tasks and showcase robustness through endurance tests. Our model was trained using an Nvidia RTX 3090 GPU and PyTorch 1.11.
5.1 Point Cloud Registration
We build upon the model architecture, training strategies, evaluation metrics, and hyperparameters established in DCP [DCP]  [^1] https://github.com/WangYueFt/dcp. We introduce the Mahalanobis version of DCP (MDCP), detailed in Section 4.1. MDCP comes in two variants: v1 without a transformer and v2 with a transformer. We further extend our approach to DeepUME [DeepUME] by proposing the Mahalanobis variant, MDeepUME. For a fair comparison, we benchmark geodesic versions of these models using vectorized Floyd-Warshall on Euclidean k-NN graphs.
Dataset. To facilitate benchmarking, we employ publicly available datasets: ModelNet40 [modelnet], FAUST [FAUST], and Stanford3D (S-3D) [^2] http:// graphics.stanford.edu/data/3Dscanrep/ . Notably, the latter two datasets are exclusively employed for testing purposes. To ensure methodological consistency, we adhere to the official training/testing splits and settings as stipulated in the original works of DCP [DCP] and DeepUME [DeepUME]. Among these datasets, our approach is rigorously evaluated.
We evaluate using root mean squared error (RMSE) between ground truth and predicted transformations (lower is better). Angular errors are reported in degrees. These metrics comprehensively assess our approach’s performance.
5.2 Comparison with State-of-the-art Methods
Table 1: Comparison on ModelNet40 [modelnet] for unseen data, unseen category, and Gaussian noise. MDCP-(v1,v2) and MDeepUME outperform baselines (bold best, underline second best). Lower is better. $\dagger$ denotes reproduced results, -Geo refers to Floyd-Warshall geodesic-graph version.
Unseen data Unseen Category Unseen data + Noise
RMSE(R) RMSE(t) RMSE(R) RMSE(t) RMSE(R) RMSE(t)
ICP [ICP] 29.9148 0.2909 29.8764 0.2933 29.7080 0.2906
GO-ICP [goICP] 11.8523 0.0257 13.8657 0.0226 11.4535 0.0231
FGR [FGR] 9.3628 0.0139 9.8490 0.0135 24.6515 0.1090
PointNetLK [pointnetLK] 15.0954 0.0221 17.5021 0.0280 16.0049 0.0216
[…]
MDCP-v1 (Ours) 1.6614 0.0004 1.9679 0.0004 1.5889 0.0004
[…]
MDCP-v2 (Ours) 0.9468 0.0004 1.3849 0.0049 0.8481 0.0043
DeepUME [DeepUME] 0.0062 0 0.0183 0.00019 2.5263 0.0006
[…]
MDeepUME 0.0023 0 0.0098 0.00009 1.9946 0.0006
ModelNet40 Results. Table 1 compare our methods (MDCP, MDeepUME) with baselines and their geodesic versions. We outperform all methods across various settings: Unseen data: Train-test split from ModelNet40 [modelnet]; Unseen category: First 20 classes for training, remaining for testing; Gaussian Noise: $\mathcal{N}(0,0.01)$ clipped to [-0.05, 0.05]. The point density is 1024 in all settings. Baseline evaluations are from [DCP] (highly comparable results achieved with minimal standard deviation during our reproduction attempts). For each case, a rigid transformation is applied along each axis (rotation: $[0^{\circ},45^{\circ}]$, translation: [-0.5, 0.5]). The network receives both the original and transformed point cloud as input, evaluated against ground truth using MDCP and MDeepUME.
Figure 3: Robustness evaluation on unseen data (FAUST [FAUST] hand), low-density source (airplane), and low-density target (chair) point clouds (trained on ModelNet40 [modelnet]). Red boxes highlights vulnerabilities in DCP-v1 [DCP], while green boxes highlights MDCP-v1’s efficacy.
Table 2: Evaluation of few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] (OBJ ONLY, OBJ+BG, PB75 splits). Our DCGNN in MDCP-(v1,v2) outperforms baselines across all settings (bold denotes best accuracy). We report mean accuracy (%) and standard deviation from 50 runs.
ModelNet40 [modelnet]
5 way 10 way
Method 10 shot 20 shot 10 shot 20 shot
DCP-v1 [DCP] 67.6 $\pm$6.91 69 $\pm$7.87 50.5 $\pm$7.74 57.1 $\pm$5.01
MDCP-v1 (Ours) 75.2 $\pm$7.3 80 $\pm$7.74 71.0 $\pm$4.16 70.9 $\pm$5.72
Increment in % 11.24 15.94 40.59 24.17
DCP-v2 [DCP] 76.8 $\pm$4.01 79.8 $\pm$6.17 60.8 $\pm$4.93 69.15 $\pm$5.7
MDCP-v2 (Ours) 82.5 $\pm$6.96 81.8 $\pm$7.78 67.9 $\pm$5.43 72.25 $\pm$6.33
Increment in % 7.42 2.51 11.68 4.48
ScanObjectNN OBJ ONLY [SONN]
DCP-v1 [DCP] 51.7 $\pm$8.08 54.8 $\pm$8.82 32.15 $\pm$3.56 34.55 $\pm$3.17
MDCP-v1 (Ours) 54.7 $\pm$7.47 57.1 $\pm$6.51 39.0 $\pm$4.88 41.2 $\pm$5.72
Increment in % 5.80 4.20 21.31 19.25
DCP-v2 [DCP] 55.1 $\pm$9.57 56.2 $\pm$6.8 38.45 $\pm$5.08 42.05 $\pm$4.93
MDCP-v2 (Ours) 55.4 $\pm$12.11 62.4 $\pm$7.91 42.85 $\pm$3.37 46.15 $\pm$2.96
Increment in % 0.54 11.03 11.44 9.75
ScanObjectNN OBJ + BG [SONN]
DCP-v1 [DCP] 50.3 $\pm$5.25 48.2 $\pm$7.35 29.45 $\pm$5.34 32.15 $\pm$2.9
MDCP-v1 (Ours) 62.2 $\pm$6.73 60.2 $\pm$7.54 37.25 $\pm$3.26 41.45 $\pm$3.21
Increment in % 23.66 24.90 26.49 28.93
DCP-v2 [DCP] 52.3 $\pm$6.92 55 $\pm$8.16 33.8 $\pm$4.31 36.25 $\pm$4.57
MDCP-v2 (Ours) 58.8 $\pm$8.07 62.8 $\pm$6.22 41.9 $\pm$6.2 44.6 $\pm$4.1
Increment in % 12.43 14.18 23.96 23.03
ScanObjectNN PB75 [SONN]
DCP-v1 [DCP] 46.1 $\pm$8.31 45.2 $\pm$9.37 25.2 $\pm$3.78 28.45 $\pm$4.5
MDCP-v1 (Ours) 54.8 $\pm$7.35 51.5 $\pm$7.08 34.05 $\pm$4.2 37.65 $\pm$4.69
Increment in % 18.87 13.94 35.12 32.34
DCP-v2 [DCP] 49.6 $\pm$6.64 49.6 $\pm$8.7 27.85 $\pm$3.75 32.35 $\pm$4.03
MDCP-v2 (Ours) 54.7 $\pm$5.89 53.2 $\pm$9.46 34.45 $\pm$3.48 41.15 $\pm$4.87
Increment in % 10.28 7.26 23.70 27.20
Few-shot Classification. Table 2 evaluates the discriminative power of Mahalanobis distance. We compare features learned by DCP-v1/v2 [DCP] and MDCP-v1/v2 in few-shot classification tasks using the $k$-way, $m$-shot setting [CrossPoint]. Models are trained for registration on ModelNet40 [modelnet] and then tested on few-shot classification for ModelNet40 and three ScanObjectNN [SONN] splits (OBJ ONLY, OBJ+BG, PB75).
MDCP consistently outperforms DCP across all settings. Summarizing Table 2, MDCP-(v1,v2) achieves average accuracy improvement of $(22.98,6.52)$ on ModelNet40. For ScanObjectNN, improvements are: OBJ ONLY: $(12.68,8.19)$; OBJ+BG: $(25.99,18.40)$; PB75: $(25.06,17.11)$, [(a, b) denotes average accuracy increment (%) for MDCP-(v1,v2)].
These results suggest two key points: 1) v1 (without transformer) outperforms v2, indicating transformers may struggle with limited data; 2) Even on unseen ScanObjectNN data, MDCP shows consistent improvement across splits.
5.3 Endurance Test
Table 3: Quantitative comparison of MDCP-(v1,v2) robustness against original on ModelNet40 [modelnet] for low-density point cloud registration. We evaluate across various settings. MDCP consistently outperforms its original counterpart (bold denotes best). Lower is better for RMSE(t) (reported in $10^{-2}$ units). Note: -Geo refers to the vectorized Floyd-Warshall geodesic-graph version.
Target Low Density ModelNet40 [modelnet]
Unseen Data Unseen Category Noise
RMSE(R) RMSE(t) RMSE(R) RMSE(t) RMSE(R) RMSE(t)
1024 DCP-v1 [DCP] 42.0026 10.865 8.3181 1.2206 32.0181 0.9616
DCP-v1-Geo 38.1172 10.863 6.0088 1.2806 39.9996 0.9801
MDCP-v1 32.6880 10.977 5.8425 1.2096 24.9222 1.0062
DCP-v2 [DCP] 68.0209 3.6777 14.9977 2.7765 66.9599 3.0532
DCP-v2-Geo 44.6676 3.7786 14.7762 8.1678 48.2278 2.9987
MDCP-v2 34.9602 3.3871 14.3836 6.1023 25.8595 2.8308
512 DCP-v1 [DCP] 54.5579 1.6058 8.31497 1.6812 50.4602 1.4233
DCP-v1-Geo 48.1628 1.7898 8.1065 1.6711 49.7866 1.5661
MDCP-v1 42.3370 1.4918 8.4701 1.6028 38.0739 1.3168
DCP-v2 [DCP] 80.8099 8.9802 23.8370 3.7857 77.4393 8.7660
DCP-v2-Geo 77.8756 6.7752 21.8861 3.6544 48.2165 8.7765
MDCP-v2 47.3271 3.9172 28.7305 8.4551 44.3240 3.4637
Source Low Density ModelNet40 [modelnet]
1024 DCP-v1 [DCP] 43.2637 1.1146 8.304669 1.228 38.6104 1.0306
DCP-v1-Geo 36.6378 1.1123 8.0011 1.2761 39.1654 1.1756
MDCP-v1 31.3686 1.1035 8.3409 1.2247 26.9862 1.0158
DCP-v2 [DCP] 65.4777 3.1448 16.1993 1.9267 61.1659 2.6327
DCP-v2-Geo 57.1998 3.0098 14.7176 6.7861 49.1109 2.9987
MDCP-v2 38.6540 2.8759 7.4574 5.6750 27.1185 2.6114
512 DCP-v1 [DCP] 55.2426 1.5553 8.4311 1.4074 51.7135 1.3908
DCP-v1-Geo 38.1172 1.0863 6.0088 1.2806 39.9996 0.9801
MDCP-v1 42.4809 1.4700 7.8250 1.1089 39.6634 1.3272
DCP-v2 [DCP] 71.8377 7.2431 22.3028 4.8566 71.1976 6.9627
DCP-v2-Geo 44.6676 3.7698 26.8861 4.9908 65.9981 5.0098
MDCP-v2 50.1560 3.6758 29.2851 4.4566 45.7104 3.3395
Efficiency towards Point-density. Table 3 evaluates MDCP’s robustness against varying point densities, simulating real-world sensor data [survey]. One point cloud (source or target) is downsampled by half (e.g., 2048 → 1024). Evaluation metrics use the same number of points for both clouds (e.g., source with 1024 points and target with 2048 points are used to compute transformation, then applied to original points for error calculation).
Figure 3 compares DCP-v1 [DCP] and MDCP-v1. When densities are similar (man’s face), both perform well. However, MDCP-v1 demonstrates superior robustness under varying densities (chair). This highlights its effectiveness in real-world applications where source or target point clouds may have different densities due to factors like sensor capabilities or data acquisition methods (e.g., SFM-LiDAR fusion).
Note: Source point cloud is a transformed version of $g(Y)$, and the goal is to estimate $g$ for perfect alignment with target $X$.
Table 4: Quantitative comparison of proposed MDeepUME’s robustness on ModelNet40 [modelnet], FAUST [FAUST], and Stanford 3D datasets. We evaluate across various noise, benchmarked by authors in DeepUME [DeepUME]. MDeepUME consistently outperforms DeepUME (bold denotes best results). Lower is better for all metrics: RMSE(t), chamfer distance (CD) [ChamferDistance], and Hausdorff distance (HD) [HausdorffDistance] (reported in $10^{-2}$ units).
Noise-types DeepUME [DeepUME] MDeepUME
CD HD RMSE(R) RMSE(t) CD HD RMSE(R) RMSE(t)
MN 40 Bernoulli 1.113 8.6300 46.39461 1.5381 1.1120 8.6200 45.7815 1.5398
Gaussian 0.1913 1.2445 2.526296 0.0625 0.1900 1.2400 1.9946 0.0642
Sampling 0.6317 5.6678 31.39413 0.9197 0.6290 5.6620 30.7709 0.9236
Z-Intersection 1.2237 11.3676 89.48223 0.9003 1.2230 11.3670 87.4486 0.911
FAUST Bernoulli 0.2608 2.7165 11.30384 2.2294 0.2728 2.7818 10.3421 2.1964
Gaussian 0.1243 0.9342 1.819818 0.1085 0.1230 0.9300 1.7029 0.2692
Sampling 0.1026 0.9342 5.6842 1.172 0.0941 0.9320 3.7985 0.9461
Z-Intersection 0.2064 2.3368 12.3721 2.2183 0.2000 2.3350 11.9663 2.1383
S-3D Bernoulli 0.2345 10.4470 5.951346 1.2268 0.2470 10.1853 6.4044 1.2266
Gaussian 0.1068 0.9477 0.392542 0.0591 0.1040 0.9410 0.3106 0.0604
Sampling 0.0908 8.5766 6.709815 0.7145 0.0900 8.4267 6.1997 0.7143
Z-Intersection 0.1771 11.7627 5.66907 1.0417 0.1700 11.5800 4.4850 1.0428
Efficiency towards Various Noise Types. Table 4 evaluates the robustness of MDeepUME against various noise types from DeepUME [DeepUME] (Bernoulli, Gaussian, Sampling, Zero-Intersection). We benchmark on ModelNet40 [modelnet] (training and evaluation), FAUST [FAUST], and Stanford3D (testing only). Training settings follow the original paper [DeepUME] ( [^3] https://github.com/langnatalie/DeepUME).
MDeepUME outperforms DeepUME across datasets and noise types in terms of Chamfer distance (CD) [ChamferDistance], Hausdorff distance (HD) [HausdorffDistance], and root mean squared error in rotation (RMSE(R)). However, limitations are observed in translation accuracy (RMSE(t)) due to DeepUME’s inherent projection into SO(3)-invariant space, where Mahalanobis distance is less effective.
6 Conclusion
Our paper introduces Mahalanobis k-NN, a statistical framework for improving point cloud registration in learning-based methods. It addresses challenges like feature matching due to variations in point cloud densities. Mahalanobis k-NN leverages local neighborhood distribution for accurate feature extraction, outperforming methods like Flyod-Warshall. It integrates seamlessly into local graph-based point cloud analysis methods like DCP and DeepUME, achieving state-of-the-art performance on benchmark datasets. We also demonstrate that registered point cloud features possess discriminative capabilities, leading to significant improvements in few-shot classification tasks. Mahalanobis k-NN offers exceptional surface awareness and outlier robustness, benefiting tasks like segmentation, decomposition, and normal estimation. Additionally, it serves as a surficial geometry-based error metric, enhancing tasks like generation and upsampling. Its versatility makes it a crucial component in advancing point cloud processing, warranting further benchmarking to highlight its significance.
[…]